OpenStudy (loser66):

How to say that l.c.m of any number in {1,2,....n} divides n! Please, help

OpenStudy (loser66):

@rational

OpenStudy (rational):

\[\large \mathbb{lcm}\times \gcd = a\times b\]

OpenStudy (loser66):

I don't get ;)

OpenStudy (rational):

lets start with that

OpenStudy (rational):

product of lcm and gcd equals produt of numbers

OpenStudy (loser66):

I mean, if I have (2)(3)(5) , then the product divides n! But I want it in general. :) English is not my mother tongue so that I have trouble in interpreting the math problem to words

OpenStudy (rational):

and the question has to be rephrased a bit you want to prove `lcm(1,2,3...n) divides n!` right ?

OpenStudy (loser66):

any of them, like (2)(3)(5) (2)(3)(5)(7) and more... until I can factor the whole set into product of something and that something divides n!

OpenStudy (rational):

Oh ok got you :) so basically we want to prove lcm of any arbitrary # of elements from the set divides n!

OpenStudy (loser66):

l.c.m ((2)(3)(5)) l.c.m(2)(3)(5)(7)) and those l.c.m divides n!

OpenStudy (loser66):

Yes. :) I am sorry for my poooooooor explanation.

OpenStudy (anonymous):

Rational's idea will work. Try to prove: $$\text{lcm}\left(a_1,a_2,\ldots ,a_k\right)\cdot\gcd\left(a_1,a_2,\ldots ,a_k\right)=a_1\cdot a_2\cdots a_k.$$

OpenStudy (loser66):

But we don't have gcd.

OpenStudy (loser66):

We just prove l.c.m (any # of the set)| n!

OpenStudy (anonymous):

Hmm...then maybe a slight change. Try to prove that if \(c\) is any common multiple of \(a_1,a_2,\ldots ,a_k \), then \(\text{lcm}\left(a_1,a_2,\ldots ,a_k\right)\mid c.\)

OpenStudy (loser66):

Or try to prove lcm (a1, a2) = c and c \(\in\) the set?

OpenStudy (anonymous):

That might not be true. The lcm(n,n-1)=n(n-1), which is too big to be in the set.

OpenStudy (loser66):

But it |n! still, right?

OpenStudy (anonymous):

ah, thats correct.

OpenStudy (loser66):

The biggest case is n(n-1)(n-2) |n! and it is the case of 3 consecutive numbers. Other smaller one is a factor of this case. Am I right?

OpenStudy (anonymous):

Intuition says yes, but that's something you would need to prove.

OpenStudy (loser66):

:(

OpenStudy (anonymous):

For any \(a_1,a_2,\cdots,a_k\in \{1,2,\ldots ,n\}\), it turns out \(n!\) is a common multiple of \(a_1,a_2,\cdots,a_k\in \{1,2,\ldots ,n\}\). That's why I think proving the lcm divides any common multiple is a decent idea. If we had access to the gcd, Rational's idea would be best and fastest.

OpenStudy (loser66):

Please , put it in logic.

OpenStudy (anonymous):

Claim: Let \(a_1,a_2,\cdots,a_k\in \{1,2,\ldots ,n\}\), let \(l=\text{lcm}(a_1,a_2,\cdots,a_k)\), and let \(c\) be any common multiple of \(a_1,a_2,\cdots,a_k\). Then \(l\mid c\).

OpenStudy (anonymous):

If we can prove this claim, then it follows that \(n!\) is always a common multiple of \(a_1,a_2,\cdots,a_k\), so \(l\mid n!\)

OpenStudy (anonymous):

To prove the claim, we use two facts: 1) Division Algorithm, and 2) The LCM is the LEAST positive common multiple. In other words, there cannot be a positive common multiple that is smaller than the LCM.

OpenStudy (anonymous):

Alright. So let \(a_1,a_2,\cdots,a_k\in \{1,2,\ldots ,n\}\), \(l\) be their LCM, and \(c\) be any common multiple. Using the division algorithm on c and l, we know there exist integers q and r such that $$c=lq+r, \quad 0\le r<l.$$

OpenStudy (anonymous):

Rearranging this equation yields: $$c-lq=r.$$ Now, \(a_i\mid c\), and \(a_i\mid l\) (because both l and c are multiples of \(a_i\)). Thus \(a_i \mid r\) for any i. But this means \(r\) is a common multiple of \(a_1,a_2,\cdots,a_k\).

OpenStudy (anonymous):

So we now know \(r\) is a common multiple of \(a_1,a_2,\cdots,a_k\), and that \(0\le r<l\). Since l is the LEAST common multiple, \(r\) cannot be positive. Therefore \(r=0\), and we conclude that $$c=lq,$$ which implies \(l\mid c\).

OpenStudy (loser66):

Question: why do we have to go on that way while we can do something shorter to get l |c? As your logic, l =l.c.m ((a1,a2,.....,a_k)) c is any common multiplication, then \(c= \alpha l\) for some \(\alpha\) in Z

OpenStudy (anonymous):

It depends on if you are given $$c=\alpha l$$ as a definition, or if you need to prove it. I wasn't sure which you were given, so i proved it just incase.

OpenStudy (anonymous):

I guess I should have asked that before I started >.< haha

OpenStudy (loser66):

:)

OpenStudy (anonymous):

Anywhos, now that you have this fact, the problem is done, since \(n!\) will be a common multiple of any \(a_1,a_2,\cdots,a_k\) in the set.

OpenStudy (loser66):

Thank you very much. I got you. :)